Shift report 20020121 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Jan 21 2002 - 00:01:34 EST

  • Next message: Apache: "Shift report 20020121 00:00-08:00"

    ------------------------------------------------------------
    Date: Mon Jan 21 00:01:34 2002, Shift: 00:00-08:00
    Supervisor: mads, with: 
    Log book#12 pages 95 - 96
    
    Shift Summary:
    ------------------------------------------------------------
    There was a long store that lasted from about 1720, and is stil running (2355). Took dataruns 6713 and 6714.
    
    Finisshed runs for MRS at 2A, and D5 magnet was set to 6A.
    
    There was a problem where the DAQ could not connect to FB3.
    After consulting JH we decided not to take data from the drift chambers. I later (2330) discovered that the reason for this was that the FB3 switch was turned of on the remote drive panel. I turned this on.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 21 2002 - 00:02:31 EST